OpenStudy (anonymous):

State all the solutions to |x-1| = 5x+2

myininaya (myininaya):

Do you know how to solve |x|=5? You have two equations to solve x=5 or -x=5 If you had |x-1|=5? You still have two equations to solve x-1=5 or -(x-1)=5 If you have |x-1|=5x+2, you still have two equations to solve. Can you tell me what two equations we need to solve? (When we are done we need to check our answer because 5x+2 isn't always positive)

OpenStudy (anonymous):

Well the question asks me "Which are the solutions to |x-1|=5x+2? Check all that apply." Then it gives me the four answer choices being A. 1/6 B. 3/4, 1/6 C.3/4 and D. -1/6? I believe I have to solve both equations just that, I'm confused where to start at.

myininaya (myininaya):

Well I was trying to get you to provide me with the two equations based off the examples I have given.

myininaya (myininaya):

Do you not understand the examples?

OpenStudy (anonymous):

No I don't understand which ones your talking about.

myininaya (myininaya):

how about some more examples then. If you have |2x-5|=4, then you have to solve 2x-5=4 and also have to solve -(2x-5)=4. because |2x-5|=2x-5 or it could be -(2x-5) depending on the values of x If you had |5x+1|=2x+3, then you have to solve 5x+1=2x+3 -(5x+1)=2x+3

myininaya (myininaya):

So if you have |x-1|=5x+2 what two equations do we need to solve?

OpenStudy (anonymous):

x-1=5x+2 and -x-1=5x+2

myininaya (myininaya):

yes we have to solve x-1=5x+2 but the other equation you have written is a little off I think you meant -(x-1)=5x+2 So we have to solve both x-1=5x+2 and -(x-1)=5x+2

myininaya (myininaya):

Do you know how to solve an equation like x-1=5x+2?

OpenStudy (anonymous):

I do but I keep getting a answer that's not on the answer choices.

myininaya (myininaya):

x-1=5x+2 So what did you do first here?

OpenStudy (anonymous):

Add 1 to the other side right?

myininaya (myininaya):

great x=5x+3

myininaya (myininaya):

then

OpenStudy (anonymous):

This is where I get confused when you subtract the five is there a one next to the x making it -4?

myininaya (myininaya):

right so you have -4x=3

OpenStudy (anonymous):

Ohh okay so is it -3/4 ?

myininaya (myininaya):

Yes but when you plug that back in will it work?

myininaya (myininaya):

And I'm talking about into original problem |x-1|=5x+2

OpenStudy (anonymous):

No it doesn't

myininaya (myininaya):

ok so we don't have a solution from that equation what about the other equation -(x-1)=5x+2

myininaya (myininaya):

First of all, don't let that negative mess you up, just go ahead and distribute (or multiply it to other side) You can write it like this -x+1=5x+2

myininaya (myininaya):

let me know what you get for x and i will tell you if you did it right

OpenStudy (anonymous):

Ohhh okay that's why I got confused I was working ahead and the negative confused me sorry hold on I'll type the answer as soon as I get it.

OpenStudy (anonymous):

I got -1/6

myininaya (myininaya):

great does it work?

OpenStudy (anonymous):

Yes I believe so

myininaya (myininaya):

Great! There is only one solution to this particular problem, being the x=-1/6

OpenStudy (anonymous):

Thank you so much can you help me with other problems too?

myininaya (myininaya):

i might can get you started but i will be running away from here soon.

OpenStudy (anonymous):

thank you what time do you usually get on here?

myininaya (myininaya):

I don't really have a set time.

myininaya (myininaya):

I come on when I can.

OpenStudy (anonymous):

The ideal length of a particular metal rod is 20.5 cm. The measured length can ideal from the ideal length 0.045. What is the range of the acceptable lengths of the rod?

OpenStudy (anonymous):

Its just asking me to write a equation.

myininaya (myininaya):

I don't understand the second sentence.

OpenStudy (anonymous):

sorry let me restate : The measured length may vary from the ideal length by the most 0.045.

myininaya (myininaya):

so do you agree that we are trying to develop an equation or inequality based off this interval: